A physical balance has its arms of unequal length. A body weighs 18 kg if kept in one pan and 8 kg if kept in the other pan. What is the true weight of the body ?

Grade:11

2 Answers

Arun
5 years ago
Dear Swastik
 
True weight of body should be = sqrt (18 * 8)
 
 = sqrt (144)
 
 = 12 kg
 
Please let me know if this is correct or not.
 
Regards
Arun
Manajs
3 years ago

Let mass of the body be m. Lengths of the beam on either side of the fulcrum of the balance be x and y.

In the first case, the body is placed on x side of the balance. Weighing block of mass 18kg is placed on the y side of the balance.

Moments of the forces about the balance are equal. Hence m.x=18.y.

In the next case 8.x=m.y

Dividing the two, we get m/8=18/m

Or, m=12kg.

Weight of the body is 12×9.8=117.6N

 
2.6K views
View Upvoters
7